Equipment acquired at a cost of $82,000 has an estimated residual value of $5,000 and an estimated useful life of 10 years. It was placed in service on April 1 of the current fiscal year, which ends on December 31.

If necessary, round your answers to the nearest cent.

a. Determine the depreciation for the current fiscal year and for the following fiscal year by the straight-line method.

b. Determine the depreciation for the current fiscal year and for the following fiscal year by the double-declining-balance method.
Answer #1

Straight line depreciation = (Cost - Salvage value)/Useful life

Current year :-

= (82,000-5,000)/10 * 9/12

= 5775

Following year :-

= 7700

Double declining rate = 200/useful life = 200/10 = 20%

Current year :-

= 82,000*10%*9/12

= 12,300

Following year

= (82,000-12,300) *25%

= 17,425